# AnyPlace: Learning Generalized Object Placement for Robot Manipulation

### v1

@@ III-B Fine-grained placement pose prediction
$\mathbf{R}\sim\mathcal{U}(\mathrm{SO}(3)),\quad\mathbf{t}\sim\mathcal{U}(\text{bbox}(\mathcal{P}_{b\_crop})).$ (3)
[…]
During training, we perform 5 denoising steps. Instead of incrementally adding Gaussian noise to the input during the forward process as is common practice [13], we manually define the noise added at each timestep. In our case, the noise is the relative transformation that the model predicts. Specifically, the intermediate ground truth relative transformations ${T}_{n,GT}^{(t)}$ are generated by linearly interpolating the translation and using spherical linear interpolation (SLERP) to sample rotations between the object’s initial and final placement poses. During inference, we generate diverse placement poses by sampling the diffusion model multiple times, each time starting with randomly transformed initial object pointclouds $\mathcal{P}_{\text{c}}^{\text{(0)}}$. We perform a larger number of denoising steps at test time, by repeating the last denoising step more times for a total of 50 denoising steps.
At each diffusion step during training, given the ground truth ${T}_{n,GT}^{(t)}$ and the predicted ${T}_{n}^{(t)}$, we use the L1 distance for the translation loss $L_{translation}$. For rotation, we measure the geodesic distance between the ground truth and predicted rotations. Additionally, we apply Chamfer loss between the pointcloud transformed by the ground truth pose and by the predicted pose. The total loss is the summation of these losses:
$\mathcal{L}_{total}=\mathcal{L}_{translation}+\mathcal{L}_{rotation}+\mathcal{L}_{chamfer}$ (4)
Figure 4: Generated objects in the synthetic dataset. We use Blender to procedurally generate various categories of objects for three placement tasks: insertion, stacking, and hanging.
III-C Robot pick and place execution
After determining the placement poses, we implement a pick-and-place pipeline to manipulate the object and position it accurately at the target pose. Specifically, we utilize AnyGrasp [7] to find viable grasps for the target object $C$ and employ cuRobo [29] as the motion planner to perform collision-free placement. We perform rejection sampling on $(T_{\text{place}},T_{\text{pick}})$ pairs to identify valid grasps for the specific placement pose predicted by our model that can be executed by the robot.
Grasp detection begins by extracting the target object pointclouds $\mathcal{P}_{\text{c}}$ from an RGBD image. AnyGrasp [7] then processes the resulting pointclouds to identify the optimal grasp candidates sorted by confidence. To pick up the object, the gripper is first moved to a pre-grasp pose, positioned 10 centimeters away from the target along the gripper’s z-axis. The gripper then approaches the target in a straight line while maintaining its orientation. Similarly, during placement, the robot first moves to a pre-place pose, followed by a final approach without altering the gripper orientation. The distance from the pre-place pose to the final placement pose is adjusted according to the object’s size to avoid collisions during the transition to the pre-place position. With the waypoints and end-effector orientation constraints defined, we use cuRobo to generate the complete motion plan for robot pick and place.
Figure 5: Robot performing various placement tasks in simulation. In the simulation, given a set of predicted placement poses and grasping poses, we leverage cuRobo for motion planning and execute all trajectories simultaneously in IsaacLab.
IV Synthetic Dataset Generation
Our aim in building the synthetic dataset is to capture a broad range of local placement arrangements. Existing models use a few very specific tasks (e.g. inserting a book into a bookshelf) to simply evaluate how well their model works given placement data for such task for training [33, 12, 32].
Our goal on the other hand is not only to use the dataset to evaluate the proposed model, but to build towards representing a broad range of types of placements (stacking, hanging, inserting) as shown in Figure 4. The local nature of our pose-prediction model makes this task much easier and enables us to build a dataset that can generalize to a broad range of real-world placement tasks.
The data generation pipeline consists of two main components: object generation and placement pose generation. Specifically, we use Blender to procedurally generate 3D objects, such as pegs, holes, cups, racks, beakers, vials, and vial holders. Object parameters—including height, width, length, and number of edges—are randomized to increase variability. For racks and vial plates, we also randomize the number of poles and holes. Additionally, random scaling is applied along the x, y, and z axes to further enhance diversity.
[…]
IV-A Training details
We train each model independently on subsets of the full dataset, split based on the type of placement configurations, as well as on the combined dataset (referred to as the ”multitask” variant). While similar methods are typically trained for a single task (e.g., [28]), our goal is to build a single model capable of performing a diverse set of placement tasks. All single-task models are trained for three days, while multitask models are trained for five days on a single NVIDIA A100 GPU. More details on training parameters are in the Appendix.
V Experimental Evaluation
TABLE I: Success rate (%) on synthetic dataset. We evaluate the baseline NSM, RPDiff, AnyPlace-EBM, and our method, AnyPlace, on the success rates of four pick-and-place tasks using objects from the synthetic dataset.
Methods Object Stacking Peg Insertion Cup Hang Vial Insertion
(single-mode) (single-mode) (multi-mode) (multi-mode)
Single task NSM $76.57$ $7.63$ $35.54$ $18.70$
[…]
Multi-task NSM $77.55$ $7.69$ $35.22$ $9.87$
[…]
AnyPlace-EBM $78.95$ $10.75$ $90.87$ $57.24$
[…]
We aim to evaluate three key aspects of our approach: success rate, coverage, and precision. With the VLM-based placement location prediction, our method effectively captures possible placement locations, achieving higher coverage rates. By narrowing the focus to minimal regions for placement pose prediction, we enable the model to achieve greater precision, which is essential for handling challenging placement tasks in real-world scenarios. All models are trained exclusively on our synthetic dataset and evaluated on novel objects in a zero-shot setting. We conduct evaluations against three baseline models on different placement tasks in both simulation and real-world settings. Simulation enables extensive testing of our approach by leveraging rejection sampling across multiple parallel environments. In real-world experiments, we highlight the effectiveness and flexibility of our VLM-based placement location prediction in handling complex environments. Additionally, we demonstrate the robustness of our method in handling pointcloud noise and generalizing to unseen objects, despite being trained on general synthetic objects in simulation and encountering novel global configurations for similar types of local placements.
Evaluation metrics. We use three metrics to evaluate model performance: success rate, coverage, and precision. Specifically, we define a placement as successful if the robot places the object at the correct location and it remains stable after release. The success rate is then calculated as the number of successful placements divided by the total number of trials. To better understand the diversity of multimodal outputs in placement prediction, we evaluate coverage, defined as the number of distinct predicted placement locations relative to the total number of possible locations. Finally, for fine placement tasks, to evaluate the precision, we measure the error between the ground truth pose and the predicted pose in terms of both distance and angle.
Baselines. We compare our approach with three baselines: NSM [1], RPDiff [28], and an energy-based model (EBM) integrated with our high-level placement location prediction module. Specifically, NSM shares the same self-attention and cross-attention pointcloud encoder, paired with a simple regression decoder. For RPDiff, since the low-level pose-prediction module in our diffusion-based approach shares the same structure, this allows us to directly examine the effects of the high-level module we propose. We do not utilize a learned classifier on top of the pose prediction model that is present in RPDiff. Such a model can be applied on top of any of the methods we evaluate here, including the ones we propose. Not having it also allows us to evaluate the number of samples needed to achieve particular coverage of possible placement locations in the scene. To evaluate the effectiveness of the diffusion decoder in generating multimodal outputs, we build an energy-based model, AnyPlace-EBM, for comparison. Inspired by Implicit-PDF [20], this model uses the same encoder as ours, but for the decoder, instead of explicitly predicting the placement pose, it includes two separate branches: one for placement location prediction and another for predicting the placement rotation energy. During training, we encourage all placement rotations in SO(3) that result in stable placements to have low energy. During inference, we randomly sample thousands of rotations and select the one with the lowest energy as the final placement rotation.
V-A Placement Success: Single & Multi-Modal
For evaluating all the methods in simulation, we utilize the pick-and-place execution pipeline we constructed (subsection III-C), together with simulation in IsaacLab [19]. The pipeline itself can be used as a standalone module. It makes no assumptions about the placement prediction method and is not specifically tied to other modules in our approach. This enables it to be used as a general system for evaluating placement pose prediction models. In Figure 5 we show the system being used to perform insertion, hanging, and placement tasks. Combined with the synthetic dataset we create, this gives us a complete system for comparing different models and getting systematic results of success rate, mode coverage, and precision. Both motion planning and the pick-and-place simulation itself are GPU parallelizable, making evaluation of new models even easier.
We evaluate single-mode and multimodal placement across a set of 4 tasks. Object stacking and peg insertion tasks are Single-model placement since only one solution is feasible. While cup hang and vial insersion are multi-modal in nature due to the existence of multiple placement locations. We show a summary of placement experiments in these simulated domains in Table I.
We first examine single-mode tasks. For the simple stacking task, where high precision in placement poses is not required, AnyPlace and baseline models, including RPDiff and AnyPlace-EBM, achieve a similar success rate of around 80%. For the peg-in-hole task, which requires high precision in placement pose prediction, AnyPlace surpasses the baseline models by a large margin.
Multi-mode tasks are where we expect to get full benefit from our approach. In the hanging task the tolerance for placing a cup on the rack is relatively high. AnyPlace achieves a 94% success rate in both single-task and multi-task settings, while RPDiff and AnyPlace-EBM perform slightly worse in comparison. Notably, NSM has the lowest success rate across all multi-mode tasks due to its inherent limitations as a regression model. In the more challenging vial insertion task, AnyPlace achieves the highest success rate of 92.74%, while the success rates of all baseline models drop significantly, with RPDiff reaching only 16.5%.
This demonstrates that relying on high-level VLM to propose possible placement locations and focusing solely on the local region for placement prediction, simplifies the task for low-level pose prediction models and enables them to better capture fine-grained pointcloud features for high-precision placements. Additionally, despite using the same high-level placement location prediction, the energy-based model suffers a performance drop of 27% and 18% in single-task and multi-task settings, respectively, compared to the diffusion-based AnyPlace. This highlights that the iterative denoising procedure in diffusion models is more effective for high-precision placement prediction.
Figure 6: Comparison of vial insertion coverage across different models. We show how the coverage ratio progresses with an increasing number of placement trials. AnyPlace clearly outperforms all baseline models as it rapidly reaches its maximum coverage.
V-B Placement Coverage: Multi-Modal Insertion & Hanging
Now we aim to investigate how the coverage changes with the increase in the number of samples taken from each model. With how NSM and RPDiff are designed, we can only take independent samples from the model until as many possible placement modes are covered. In the case of AnyPlace and AnyPlace-EBM, the high-level VLM module predicts possible placement locations, so we can perform sampling for each mode. We perform an equal number of samples in each and compare performance for the same total number of samples for each model. All models used in this evaluation are trained on the multitask dataset.
In Figure 6 we show results for the vial insertion task. AnyPlace achieves close to its maximum performance after with a single sample at each mode, outperforming other models with 100 samples. This is due to its high placement success rate and the VLM’s strong reasoning ability in identifying various possible placement locations. A similar trend is observed for the AnyPlace-EBM models; however, due to a lower placement success rate, the model plateaus at around 73%. In contrast, RPDiff, despite being a diffusion-based model that should capture multimodal outputs, fails in this regard, with coverage fluctuating below 10%. Even the NSM regression model outperforms it in this case, but still stays far below either of the methods utilizing the high-level VLM module.
For hanging, to further evaluate the effect of different objects on predicted placement coverage, we aim to assess how these models generalize to racks with different sizes, geometries, and spacing between the sticks. Specifically, we generate new racks with different physical parameters compared to those used during training. The success rates are shown in Table II, and the coverage is plotted in Figure 7. Similar to the vial insertion task, AnyPlace outperforms other baselines, consistently getting higher coverage with fewer samples and eventually reaching 100% coverage – made possible with the exceptional performance of the VLM module in finding the placement modes. RPDiff performs better in this case, but still struggles to cover all placement locations, with coverage saturating around 90%. This illustrates that the original RPDiff does not generalize well to objects with different geometries and sizes compared to the training set. In contrast, AnyPlace, aided by high-level placement prediction, allows the low-level prediction model to focus on learning placements based on local regions, thereby enabling strong generalization across varying object sizes and shapes.
Figure 7: Comparison of hanging coverage across different models. AnyPlace achieves nearly perfect coverage by leveraging the VLM’s ability to propose possible placement locations, whereas other baseline models fail to capture all placement modes.
V-C Placement Precision: Fine-Grained Insertion
Figure 8: Distance and angle errors on insertion tasks. We plot histograms of the translation and rotation error in pose prediction for each approach. We base the error on the position and rotation to the closest viable placement pose. (MT) denotes a model trained on multi-task data.
The final evaluation we conduct in simulation focuses on assessing the precision of each approach. Precision is a critical factor in the successful completion of many placement tasks. Our goal is to determine whether providing models with a much smaller input region can enhance the precision of placement pose estimation. To achieve this, we evaluate the output of the prediction model directly rather than analyzing the results of the placement in simulation.
In Figure 8, we show the distribution of distance errors for each approach on the insertion tasks. Within model outputs that predict approximately the correct location, AnyPlace produces smaller errors and does so more reliably than the baselines. We also present the corresponding results for rotation prediction. Since the objects being placed are symmetric around the z-axis, we exclude yaw angle prediction from this comparison. As the initial pose is randomized and approaches cannot infer this orientation, all methods produce uniformly random yaw angle values. Given the minimal difference between the top and bottom sides of the object in this task, all approaches—except NSM, which performs the worst in both position and orientation—predict correct and flipped poses with roughly equal likelihood. Both AnyPlace and the RPDiff baseline perform well in predicting the correct orientation. Notably, the AnyPlace-EBM model achieves only slightly lower performance compared to the diffusion-based RPDiff and outperforms the NSM baseline. This is significant, as it highlights the potential of non-diffusion-based models for tackling these types of problems in future research.
TABLE II: Success rate of the hanging task across different models. Using the evaluation procedure outlined in this section, we calculate the success rate of the hanging task on a rack with different object characteristics.
Methods Success Rate
NSM Multitask 37.26%
RPDiff Multitask 81.17%
AnyPlace-EBM Multitask 69.43%
[…]
Figure 9: Robot performing various placement tasks in real-world environments. In the real-world setting, with a simplified placement pipeline, the robot successfully executes a variety of placement tasks across a wide range of placement locations based on the prediction of the AnyPlace model. (A) We demonstrate peg-in-hole, cup-on-rack, and plate-on-rack tasks. (B) The robot places the bottle in different locations based on a language description. (C) The robot stacks rings of different radii. (D) The robot executes a long-horizon task where it places the lid on the pot and then places the pot on the stove.
V-D Real World Evaluation: Precise Multi-Modal Placement
TABLE III: Success rate of real robot executing vial insertion task. Our model significantly outperforms baseline models in real experiments, demonstrating its ability to generalize to unseen objects and effectively handle noisy data.
Methods Number of Success
NSM 0/10
RPDiff 0/10
AnyPlace-EBM 5/10
[…]
We also evaluate the performance of our approach on various placement configurations with different objects in the real world. Specifically, for each scene, we capture a single RGBD image with a ZED Mini camera mounted on a Franka Emika Robot arm. We use exactly the same high-level pipeline for detecting local placement locations as we do in our simulation experiments. We also utilize the same models, trained purely on our synthetic dataset, for predicting the placement poses in a zero-shot manner. We use a simplified pipeline for executing the placement, using a specific grasp and performing placement inverse kinematics instead of a full motion planner. In addition, we do not utilize rejection sampling, but directly execute each trajectory on real. Our main goal is to evaluate whether enabling the model to predict only local placements would allow our models to generalize well to unseen objects in the real world even when only trained on our synthetic dataset. We perform systematic evaluations on the vial insertion task, with multiple possible placement locations. We also execute the approach on other language-conditioned placement variants in the real world, examples of which can be seen in Figure 9. All prompts used by the VLM to predict placement locations can be found in the Appendix.
We use a standard vial rack used in chemistry labs and randomly block different holes with other vials or cover them with tape to test the flexibility of our high-level placement location prediction module. Overall, across 10 trials, the NSM and RPDiff fail to adapt to differences in this real-world task compared to the training conditions. AnyPlace, on the other hand, successfully completes placements, achieving the best performance, and inserts vials into 8 out of the 10 available empty spots in the vial rack.
V-E Generalizable & Language-Conditioned Placement: Applicability in Real-World Tasks
To evaluate the generalization of AnyPlace, we conduct experiments on a large variety of objects for different placement configurations. As shown in Figure 9 (A), our method enables the robot to successfully perform placements in different scenarios, such as inserting a peg into a hole, hanging a cup on an unseen rack, and placing a plate into different slots of a plate rack. For precise placement, we demonstrate in Figure 9 (C) that the robot can accurately stack rings of varying radii, showcasing its ability to handle fine-grained placement tasks. Beyond single-step placements, AnyPlace excels in long-horizon tasks. For instance, we instruct the robot to precisely place a lid on a pot before placing the pot on a stove. Our method ensures that placed objects are centered and properly aligned, highlighting its reliability in real-world tasks. To demonstrate the advantage of combining VLM-based placement location prediction and local placement pose prediction modules, we show that the robot is able to place a bottle at different locations on a shelf based on language descriptions, as shown in Figure 9 (B). The key advantage of AnyPlace lies in its ability to integrate VLM-based high-level placement prediction with local placement pose refinement. By focusing on local regions for placement prediction and leveraging a diverse synthetic training dataset, it allows AnyPlace not only to generalize to real-world objects with significantly different geometries but also captures complex placement configurations with greater accuracy and flexibility.
VI Limitations
Our primary focus in this work is on predicting placement poses. However, challenges remain in executing the full pick-and-place task with the same level of generality. Not every stable grasp of an object can be used to place it at a specific pose, and performing rejection sampling in real-world scenarios can be both difficult and time-consuming. Nonetheless, we believe our synthetic dataset and evaluation pipeline provide a strong foundation for advancing in this direction by enabling the generation of training data for an end-to-end pick-and-place model applicable to a wide range of real-world placement tasks.
While our approach improves precision in placement pose prediction, it is still limited by the accuracy of the point clouds it receives as input. Completing placement tasks that require high precision can be challenging with imperfect point cloud data. Recent advances in depth estimation from RGB images show promise in addressing this issue, as does continued progress in sensor quality. Another promising approach is implementing a policy that uses force/torque feedback to refine the final stage of placement. Similarly to the previous point, we believe our dataset and simulation pipeline provides a strong foundation for tackling this challenge by generating data for training reactive placement-execution policies.
The approach we propose does not include language conditioning in the low-level pose prediction model. As a result, we cannot distinguish between different types of placements at the same location. However, given the architecture of the low-level model, adding a language-conditioning input is straightforward. Our synthetic data generation pipeline is well-suited for this scenario, as it allows for the automatic addition of language conditioning to different placement types. Additionally, an LLM can be used to enhance diversity in language-conditioning data.
VII Conclusion
In this work, we presented a general pipeline for performing a wide range of object placement tasks using a robotic arm. We proposed a two-part framework consisting of a high-level module that determines potential placement locations and a low-level module that predicts fine placement poses. The core idea of our approach is to leverage a VLM to propose placement locations, allowing the low-level pose prediction model to focus only on the local region of interest in the object’s pointcloud. This effectively reduces complexity and enhances generalization. To train our model, we created a synthetic dataset containing thousands of randomly generated objects and placement poses. We demonstrated the effectiveness of the entire pipeline in both simulation and real-world experiments. In simulation, we showed that AnyPlace outperforms baseline methods in terms of success rate, coverage, and precision. We then validated its robustness and generalization in real-world settings, where, given a single RGB-D image, AnyPlace predicts diverse placement configurations in a zero-shot manner and successfully generalizes to unseen objects.
Acknowledgments
We would like to acknowledge the part of the funding provided by the University of Toronto’s Acceleration Consortium from the Canada First Research Excellence Fund, grant number CFREF-2022-00042. We would also like to thank Dr. Anders G. Frøseth, the Vector Institute, the Acceleration Consortium, Natural Resources Canada, and the Canada 150 Research Chairs program for their generous support.
Appendix
VII-A Language Prompts for Molmo and Predicted Placement Visualization
In Figure 10 and Figure 11, we show additional language prompts used by the VLM to predict placement locations in both real-world and simulation experiments. Based on our observations, the Molmo VLM accurately identifies the correct placement locations in both real-world and simulated images based on the language input. Even when the predicted location is not perfectly centered, our low-level pose prediction model still works. For instance, when predicting the placement position of the bottle on the top shelf, the location may be at the edge. However, our model focuses on the local region centered at the proposed placement location and provides a pose that allows the robot to execute the task successfully.
VII-B Additional Details on Model Training and Dataset
In Table IV, we list all the parameters used for training the AnyPlace low-level pose prediction models. Table V presents statistics on the number of placements generated using our synthetic dataset generation pipeline. Notably, by focusing only on the region of interest for placement pose prediction, AnyPlace models perform well across different placement tasks and generalize to unseen objects. They are trained with fewer than 2,000 samples per task, demonstrating the effectiveness of our design.
TABLE IV: Parameters for Model Training
Parameter Value
Diffusion steps 5
Number of training iteration 500k
Batch size 48
Optimizer AdamW [18]
Learning rate $1\times 10^{-4}$
Learning rate schedule Linear warmup and cosine decay
Warmup epochs 50
Weight decay 0.1
Optimizer momentum $\beta_{1}=0.9$, $\beta_{2}=0.95$
TABLE V: Dataset Size for Each Placement Task
Placement Task Number of Placements
Hanging 1,767
Stacking 1,696
Vial Insertion 1,107
Other Insertion 800
VII-C Additional Details on Model Architecture
For our diffusion-based placement prediction model, we leverage a transformer architecture. We use self-attention layers to extract point cloud features for individual objects and cross-attention to learn joint features for placement pose prediction. Details of the model architecture are listed in Table VI.
Figure 10: Additional Language Prompts and VLM Output Visualization in the Real-World Evaluation.
Figure 11: Language Prompts and VLM Output Visualization in the Simulation Evaluation. In our simulation experiments, we use the VLM to generate placement locations on RGB images from the simulator. Based on this proposed location, we crop the point clouds and input them into our placement pose prediction model.
TABLE VI: Summary of Model Architecture

### v2

@@ B.1 Model Architecture
$T_{\text{init}}=(\mathbf{R},\mathbf{t}),\quad\mathbf{R}\sim\mathcal{U}(\mathrm{SO}(3)),\quad\mathbf{t}\sim\mathcal{U}(\text{bbox}(\mathcal{P}_{b\_crop})).$ (2)
As shown in Figure 2, at each denoising timestep $t$, $\mathcal{P}_{\text{c}}^{\text{(t)}}$ and $\mathcal{P}_{\text{b\_crop}}$ are input into the encoder. Specifically, both point clouds are first downsampled to $1024\text{\,}\mathrm{p}$oints using Farthest Point Sampling (FPS) and normalized to the size of a unit cube. The downsampled point clouds are passed through a linear layer to extract latent features, which are subsequently concatenated with a one-hot vector used to identify the corresponding point cloud. These combined features are then processed by the Transformer encoder [22, 12], where self-attention layers are applied to effectively extract features from the point clouds. We then leverage cross-attention and pooling layers to further aggregate these features, producing a unified feature representation that captures the spatial relationship between the two objects. In the decoder, we first obtain the sinusoidal positional embedding of the diffusion timestep $t$. Finally, the joint point cloud feature representation, along with the encoded timestep, is fed into MLP layers to predict the relative transformation ${T}_{n}^{(t)}$ consisting of a rotation $\mathbf{R}\in\text{SO}(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$ for refining the object’s pose. The target object point clouds are then transformed accordingly before proceeding to the next denoising step as $\mathcal{P}_{\text{c}}^{\text{(t-1)}}={T}_{n}^{(t)}\mathcal{P}_{\text{c}}^{\text{(t)}}$. The full transformation $T_{n}$ taking the object from its initial location to the placement pose is the product of all the incremental transformations predicted through the diffusion steps: $T_{n}=\prod_{t=1}^{t_{\max}}T_{n}^{(t)}$.
We do not utilize a learned classifier on top of the pose prediction model that is present in RPDiff. Such a model can be applied on top of any of the methods we evaluate here, including the ones we propose. Not having it also allows us to evaluate the number of samples needed to achieve a particular coverage of possible placement locations in the scene.
Table 3: Summary of Model Architecture
@@
B.2 Model Training
During training of our low-level diffusion pose prediction model, we perform 5 denoising steps. Instead of incrementally adding Gaussian noise to the input during the forward process, as is common practice [28], we manually define the noise added at each timestep. In our case, the noise is the relative transformation that the model predicts. Specifically, the intermediate ground truth relative transformations ${T}_{n,GT}^{(t)}$ are generated by linearly interpolating the translation and using spherical linear interpolation (SLERP) to sample rotations between the object’s initial and final placement poses. During inference, we generate diverse placement poses by sampling the diffusion model multiple times, each time starting with randomly transformed initial object point clouds $\mathcal{P}_{\text{c}}^{\text{(0)}}$. We perform a larger number of denoising steps at test time, by repeating the last denoising step more times for a total of 50 denoising steps, similar to RPDiff [1]. All single-task models are trained for three days, while multitask models are trained for five days on a single NVIDIA A100 GPU. Other parameters can be found in Table 4.
Table 4: Parameters for Model Training
Parameter Value
Diffusion steps 5
Number of training iteration 500k
Batch size 48
Optimizer AdamW [29]
Learning rate $1\times 10^{-4}$
Learning rate schedule Linear warmup and cosine decay
Warmup epochs 50
Weight decay 0.1
Optimizer momentum $\beta_{1}=0.9$, $\beta_{2}=0.95$
B.3 Additional Details on Data Generation
For object generation in Blender, we procedurally generate 3D objects, such as pegs, holes, cups, racks, beakers, vials, and vial holders. Object parameters—including height, width, length, shapes, and number of edges—are randomized to increase variability. For racks and vial plates, we also randomize the number of poles and holes. Additionally, random scaling is applied along the x, y, and z axes. This type of programmatic object generation essentially gives us placement poses ”for free”.
This approach for object generation not only allows simple randomization of object geometries, but also supports a range of valid placements rather than a single fixed pose. For training the AnyPlace low-level pose-prediction model, only local placement data is needed, so we do not have to generate full objects with all possible placement modes, further simplifying scalability. While adding completely new object classes does require some human effort, it is limited to writing a single piece of code per class, and common geometric aspects can often be shared. Overall, building a large dataset of parametrized object classes enables programmatic creation of vast amounts of synthetic training data, supporting not only placement tasks but potentially many other manipulation problems.
To identify stable placements for objects, we use NVIDIA IsaacSim to determine object placement poses. At the start of each trial, two objects are randomly sampled and loaded into the simulation. Since all objects are procedurally generated and possible placement locations are known during generation (e.g., the center of each hole on a vial plate), the ideal object placement location for the placing object can easily be determined. This approach finds object placement locations that maximize the clearance between the objects in their final placement configurations. For object placement rotations, they are then randomly sampled along their axis of symmetry to explore various placement poses. Four cameras are set up to capture dense object point clouds and render RGBD images. This dataset covers a wide range of placement scenarios encountered in real life. Table 5 presents statistics on the number of placements generated using our synthetic dataset generation pipeline. Notably, by focusing only on the region of interest for placement pose prediction, AnyPlace models perform well across different placement tasks and generalize to unseen objects. They are trained with fewer than $2000\text{\,}\mathrm{s}$amples per task, demonstrating the effectiveness of our design.
Table 5: Dataset Size for Each Placement Task
Placement Task Number of Placements
Hanging 1,767
Stacking 1,696
Vial Insertion 1,107
Other Insertion 800
Appendix C Evaluation
C.1 Implementation of the Pick and Place Pipeline
After determining the placement poses by AnyPlace, we implement a pick-and-place pipeline to manipulate the object and position it accurately at the target pose. Specifically, grasp detection begins by extracting the target object point clouds $\mathcal{P}_{\text{c}}$ from an RGBD image. AnyGrasp [23] then processes the resulting point clouds to identify the optimal grasp candidates sorted by confidence. To pick up the object, the gripper is first moved to a pre-grasp pose, positioned 10 centimeters away from the target along the gripper’s z-axis. The gripper then approaches the target in a straight line while maintaining its orientation. Similarly, during placement, the robot first moves to a pre-place pose, followed by a final approach without altering the gripper orientation. The distance from the pre-place pose to the final placement pose is adjusted according to the object’s size to avoid collisions during the transition to the pre-place position. With the waypoints and end-effector orientation constraints defined, we use cuRobo to generate the complete motion plan for robot pick and place while accounting for the object collision model. For each evaluation scene, we perform rejection sampling by generating 100 grasps using AnyGrasp and forming (grasp, placement pose) pairs to identify valid grasps for the specific placement pose predicted by our model that can be executed by the robot.
The pipeline itself can be used as a separate module. It makes no assumptions about the placement prediction method and is not specific in any way to other modules in our approach. This enables it to be used as a general system for evaluating placement pose prediction models. In Figure 3 we show the system being used to perform insertion, hanging, and placement tasks. Combined with the synthetic dataset we create, this gives us a complete system for comparing different models and getting systematic results of success rate, mode coverage, and precision. Both motion planning and the pick-and-place simulation itself are GPU parallelizable, making the evaluation of new models even easier.
C.2 AnyPlace Energy-based Model
Inspired by Implicit-PDF [26], this model uses the same encoder as ours, but for the decoder, instead of explicitly predicting the placement pose, it includes two separate branches: one for placement location prediction and another for predicting the placement rotation energy. During training, we encourage all placement rotations in SO(3) that result in stable placements to have low energy by predicting the unnormalized log probability of the joint distribution between the latent state from the encoder and the SO3 rotation. This approach requires only minimal changes to the model. We use a negative log-likelihood loss for the rotation as shown below:
$\mathcal{L}_{energy}(h,{R}_{0})=-\log\,\frac{\exp(f(h,R_{0}))}{\sum_{i}^{N}\exp(f(h,R_{i}))}$ (3)
During training, we estimate the normalization factor by sampling. In particular, we ran the model on 4096 randomly sampled SO3 rotations. The total loss is the sum of this loss and the same L1 loss for translation:
$\mathcal{L}_{total}=\mathcal{L}_{translation}+\mathcal{L}_{energy}$ (4)
At test time, we can get the estimate of the full distribution over the rotations conditioned on the inputs to the model in the same way by randomly sampling thousands of rotations. The rotation with the lowest energy is then selected as the final placement rotation.
We found that the translation loss consistently converges faster than the energy-based one during training as a result of its smaller scale. To address this imbalance, we experimented with various ways of weighting the two loss terms when computing the total loss, which led to slight performance improvements. In contrast, for the diffusion model, losses are computed at each diffusion step, and the scales of translation and rotation errors remain more balanced, resulting in more stable training. Overall, the energy-based model offers faster inference but lower accuracy compared to the diffusion model.
Appendix D Additional Result of Placement Precision
In this section, we provide additional results on the placement precision of fine-grained insertion tasks for all models trained on single-task data. In Figure A3, we show the distribution of distance errors for each approach on the insertion tasks. For both single-task and multi-task training, we observe that AnyPlace achieves smaller errors and does so more reliably than the baselines, in terms of translation and rotation prediction.
Figure A3: Evaluation of Distance and Angle Errors in Insertion Tasks. We plot histograms of the translation and rotation error in pose prediction for each approach trained on single-task data and multi-task data. We base the error on the position and rotation to the closest viable placement pose. The rotation error is in the range $0,\pi]$. For both translation and rotation, we use 50 equally-sized bins covering the range of the variable to produce the histograms.
D.1 Additional Analysis of Real Robot Experiments
Figure A4: Object point clouds at predicted placement poses in failure cases. The point cloud of the object being placed is shown in black.
We systematically evaluate the performance of all baseline models and AnyPlace across three tasks on a real robot: vial insertion, ring hanging, and battery stacking. As shown in the Figure A5, AnyPlace successfully predicts correct and precise placement poses using partial point clouds for all tasks, due to its ability to focus on the local placement region and the robustness of its diffusion-based pose prediction model. Although AnyPlace-Energy uses the same high-level placement location module, we observe that the energy-based model often predicts incorrect poses in all tasks, both in position and orientation. Similarly, RPDiff frequently converges to the same placement pose, failing to capture diverse placement locations. It also struggles significantly with predicting accurate positions. Due to the fixed cropping size introduced in RPDiff, its generalization to various objects is limited, causing inaccurate prediction. For example, in the vial insertion and ring hanging tasks, the predicted poses often cause the placing object’s point cloud to deeply intersect with the base objects. For battery stacking, the model predicts a pose that is far from the correct placement location. Finally, NSM, which uses a simple regression decoder, is unable to predict any meaningful placement poses.
Figure A5: Visualization of Object Point Clouds at Predicted Poses from Different Models. Quantitative evaluations were conducted on three real-world tasks: vial insertion, ring hanging, and battery stacking. The figure shows examples of predicted placement poses visualized using point clouds. The point cloud of the placed object (target object) is shown in black.
To further evaluate the generalization and robustness of AnyPlace in the real world, we conducted extensive experiments across different placement scenarios. We visualize the point clouds at the predicted placement poses in Figure A6. It is clear that AnyPlace is capable of handling unseen objects and noisy point clouds. For example, in many insertion tasks, the model accurately predicts placement poses even when dealing with complex, unseen objects, such as focusing on individual fingers when hanging a ring onto a hand. This level of generalization is achieved by focusing on local regions rather than relying on global information, which may contain redundant or irrelevant details. Moreover, training the low-level model on a diverse synthetic dataset enables it to capture general concepts and representations across different placement scenarios. For example, the model was never trained with a funnel or the ring clamp, yet it successfully predicts the correct placement poses for these novel objects.
Figure A6: Object point clouds at predicted placement poses in failure cases. The point cloud of the object being placed is shown in black.